# Multiple Queries with Multiple Keys: A Precise Prompt Matching Paradigm for Prompt-based Continual Learning

### v2

@@ 4.1 The Design Goal
For question 1, we employ a ViT with prompts as the query function. A ViT with prompts not only incorporates ViT’s generalized knowledge but also task-specific fine-grained knowledge, enabling more precise feature extraction for queries. For question 2, we introduce class-level learnable parameters as keys. Since samples within a class tend to be similar, class-level keys can represent the feature distribution of a task’s samples better. With queries and keys in place, the model can evaluate the similarity between the query and keys, selecting the most consistent prompt for the sample to make the final prediction.
4.2 Multiple Queries Matching Multiple Keys
@@
Given a sample $\mathbf{x}$, we use the $t$-th e-prompt and g-prompt in combination with ViT for feature extraction to obtain $f_{r}(P_{g};P_{t};\mathbf{x}_{e})[0]$, which serves as the $t$-th query $Q_{t}\in\mathbb{R}^{D}$. The queries generated by using different e-prompts form a query pool $\mathbf{Q}=\{Q_{1},Q_{2},\cdots,Q_{M}\}$. We refer to task-level queries for breadth search as MQ. Next, $Q_{t}$ is matched with $\mathbf{k}_{tj}$ by calculating the cosine similarity $\cos\langle Q_{t},\mathbf{k}_{tj}\rangle$. At this point, each prompt has matching scores for all the categories within the task. We need to select the top-$K$ categories with the highest score within a task for aggregation by:
$S^{t}=\underset{\{c_{m}\}_{m=1}^{K}\subseteq[1,|\mathcal{Y}^{t}|]}{\max}\sum_{m=1}^{K}\cos\langle Q_{t},\mathbf{k}_{tc_{m}}\rangle,$ (2)
where $S^{t}$ is the aggregated matching score between the query and the $t$-th prompt. Finally, we select the highest matching score, and get the corresponding index as follows:
$I=\underset{t\subseteq[1,M]}{\operatorname{argmax}}\ S^{t}.$ (3)
@@
4.3 Optimization Objective
The overall optimization objective is divided into two parts: one is the learning of the prompts, and the other is the learning of the keys. The prompts are learned by optimizing the cross-entropy loss between the predicted results and the ground truth by:
$\mathcal{L}_{\text{prompts}}=\text{CE}(W^{T}f_{r}(P_{g};P_{t};\mathbf{x}_{e})[0],y),$ (4)
@@ 4.3 Optimization Objective
Since queries have already been dispersed through cross-entropy in the classification objective, the objective for the keys is solely to align with the queries. Hence, the loss for learning the keys can be written as:
$\mathcal{L}_{\text{keys}}=\sum_{i=1}^{M}\mathbb{I}(i=t)\sum_{j=1}^{|Y^{i}|}(1-\cos\langle Q_{i},\mathbf{k}_{ij}\rangle)\mathbb{I}(j=y),$ (5)
@@
where $\mathbb{I}(\cdot)$ is the indicator function. Our MK learns by aligning with the MQ that has already been dispersed through cross-entropy. In contrast, the MK in CPrompt learns by dispersing itself through cross-entropy, since SQ is fixed and not dispersed. Our MK is a component specifically adapted for MQ.
@@
As shown in Equation 5, keys corresponding to other categories and queries corresponding to other prompt are not involved in the loss calculation. This means that during training, only one query is required, implying that the MQ mechanism not only does not introduce additional computational burden, but actually requires one less query without prompt compared to SQSK in the training process. The overall loss function can be written as:
$\mathcal{L}=\mathcal{L}_{\text{prompts}}+\mathcal{L}_{\text{keys}}.$ (6)
4.4 Discussion on Inference Computational Costs
MQ needs to be executed in the inference process, so $M$ forward passes through the ViT backbone are required to obtain all queries. In contrast, SQSK only requires one query without the prompt and one prediction with the selected prompt. Therefore, the computational cost of MQMK is approximately $M/2$ times that of SQSK. The different $Q_{t}$ queries are independent of each other. Therefore, if computational resources are sufficient, the computation time of MQMK can be reduced to $1/{M}$ of the original time. The is a theoretical analysis, and the practical results, the parameters analysis and computational costs during the training phase, are discussed in Appendix B.
@@ 4.4 Discussion on Inference Computational Costs
CPrompt - - 77.14 $\pm$ 0.11 5.97 $\pm$ 0.68 74.79 $\pm$ 0.28 7.34 $\pm$ 0.65
[…]
Method $A_{T}$ ($\uparrow$) $F_{T}$ (($\downarrow$( ↓)
@@
[…]
5 Experiments
5.1 Evaluation Benchmarks
Dataset. We shuffle the classes and perform task splitting on three widely used visual datasets for prompt-based continual learning: CIFAR-100 (Krizhevsky et al., 2009), ImageNet-R (Hendrycks et al., 2021), and DomainNet (Peng et al., 2019), in order to align with the problem setup and conduct comprehensive experiments. We divide CIFAR-100 and DomainNet into 10 tasks, while ImageNet is divided into three cases: 5 tasks, 10 tasks, and 20 tasks. The dataset introduction is in Appendix F. The data distributions of DomainNet and ImageNet-R differ significantly from the ImageNet dataset used for model pre-training. This requires the model to continually learn new knowledge from these datasets, rather than relying on the knowledge learned from the pre-training dataset.
Evaluation Metrics. Average accuracy (Kim & Han, 2023; Yan et al., 2021; Li & Hoiem, 2017) and forgetting rate (Chaudhry et al., 2018; Lopez-Paz & Ranzato, 2017) are the two core metrics we use. $A_{T}$ is the average accuracy on tasks 1 to $T$ after the model has learned task $T$, and can be computed by:
$A_{T}=\frac{1}{T}\sum_{t=1}^{T}\text{Accuracy}(t,T),$ (7)
@@ 5.1 Evaluation Benchmarks
$F_{T}=\frac{1}{T}\sum_{t=1}^{T}\left(\text{Accuracy}\left(t,t\right)-\text{Accuracy}\left(t,T\right)\right).$ (8)
@@
Implementation Details. ViT-B/16 (Dosovitskiy, 2020) trained on ImageNet-1K (Russakovsky et al., 2015) is the backbone we use, so $S$ is set to 16. G-prompt is used in the first two layers of MSA, with a length of 5. The depth and length of e-prompt are discussed in Section 5.5. In all experiments, $M$ is set equal to $T$, meaning that only one prompt is learned for each task. Each prompt selects only the category with the highest similarity for aggregation which implies that $K$ is set to 1. For both SQSK and MQMK, a learning rate of 0.005, a batch size of 64, and the Adam (Kingma, 2014) optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ are used in all experiments.
Method $A_{T}$ ($\uparrow$) $F_{T}$ (($\downarrow$( ↓)
@@
[…]
Table 3: Results (%) on Split DomainNet under 10-task setting. Best results are marked in bold. All our results are over 5 trials. Except for MQMK, all the other results come from (Gao et al., 2024).
@@
Figure 4: Ablation experiments on the 10-task Split ImageNet-R. The average accuracy and matching rate change with the learning process.
Competitors and Joint Training. We compare our method with SOTA prompt-based continual learning models, including L2P (Wang et al., 2022c), DualPrompt (Wang et al., 2022b), ESN (Wang et al., 2023), CODA-P (Smith et al., 2023), EvoPrompt (Kurniawan et al., 2024) and CPrompt (Gao et al., 2024). To ensure a fair comparison, we use SQSK as a baseline in some experiments. SQSK and MQMK have identical prompts, with the only difference being the matching mechanism. SQSK can be seen as a version of DualPrompt with adjusted hyper-parameters, and it has nearly the same performance as DualPrompt reported in Tables 1, 2 and 3. Additionally, we jointly train a model using data from all tasks with linear prototypes and prompt tuning as a reference for the upper bound.
5.2 Performance Comparison
As shown in Table 1, in the three task settings on ImageNet-R, MQMK achieves SOTA performance. In the 5-task and 10-task splits, MQMK shows no significant performance gap compared to joint training. For the small task split, the continual learning performance with MQMK no longer shows a significant gap compared to traditional joint training with all data. This suggests that the performance on ImageNet-R with the 5-task and 10-task splits may have already approached its oracle. However, on the 20-task split, the improvement with MQMK is relatively small. This may be due to each query generating a subspace, and long-term continual learning leads to an increase in the number of subspaces. Different subspaces might overlap to some extent, making it difficult to perform precise query prompting. Additionally, compared to the high-performing CPrompt and EvoPrompt, MQMK has a clear advantage in terms of forgetting rate. This indicates that MQMK not only achieves higher final performance, but also experiences smaller performance degradation during the learning process. CODA-P has a low forgetting rate, yet its average accuracy is also relatively low. This suggests that CODA-P exhibits overall stability in performance, but its final performance is average.
5.3 Ablation Study
The components of our method include Multiple Queries (MQ) and Multiple Keys (MK). The corresponding components in existing methods are Single Query (SQ) and Single Key (SK). We discuss the average accuracy and matching rate under four combinations: MQMK, MQSK, SQMK, and SQSK, as illustrated in Figure 4. Using SQSK as the baseline and adding MK, $A_{T}$ improves by 2.84% and the matching rate increases by 11.32%. This demonstrates the effectiveness of setting class-level keys for each prompt. By adding MQ, $A_{T}$ decreases by 8.13% and the matching rate drops by 8.92%. This is because, under the supervision of cross-entropy, the features of samples from the same class are more clustered, while the features of samples from different classes are more dispersed. The features of multiple classes in a task have already been dispersed, and in this case, SK can no longer represent the feature distribution of all training samples in a task. When MQ is combined with MK, $A_{T}$ increases by 8.21% and the matching rate improves by 32.82%. This demonstrates that incorporating task-related knowledge significantly enhances retrieval accuracy. In fact, in query-key matching, the improvement of query quality and key quality are complementary. Enhancing both simultaneously has a significant impact on matching rate.
Number $A_{T}$ ($\uparrow$) Matching Rate (($\uparrow$( ↑)
@@ 5.3 Ablation Study
Table 4: $A_{T}$ (%) and matching rate (%) under keys of different granularities on 10-task Split Imagenet-R.
5.4 Setting Fine-grained Keys
MQSK leads to a performance decline, whereas MQMK improves performance. This highlights the critical importance of learning reliable keys for distribution representation. Based on MQ, we gradually increase the number of keys from 1 to the number of classes in each task (20 class per task in 10-task Split Imagenet-R), and we report their performance in Table 4. Both SK and MK are special cases of $N$-keys. It can be observed that as the number of keys increases, both the matching rate and accuracy improve, and the optimal performance is achieved when the key reaches the class level. Since the smallest granularity of our labels is at the class level, when the number of keys exceeds the number of classes, it becomes difficult to provide appropriate supervisory signals for the keys.
@@ 5.4 Setting Fine-grained Keys
Figure 5: Average accuracy of MQMK and SQSK on the 10-task Split ImageNet-R under different depths and lengths of the e-prompt. (a) The length is fixed at 80, and the depth varies. (b) The depth is fixed at 10, and the length varies.

### v3

@@ 4.1. The Design Goal
  3. (3)
    How can we effectively establish a matching mechanism between the features of test samples and the distribution of training samples?
For question 1, we employ a ViT with prompts as the query function. A ViT with prompts not only incorporates ViT’s generalized knowledge but also task-specific fine-grained knowledge, enabling more precise feature extraction for queries. For question 2, we introduce class-level learnable parameters as keys. Since samples within a class tend to be similar, class-level keys can represent the feature distribution of a task’s samples better. For question 3, to reduce the interference introduced by queries generated from different prompts, we employ a local matching strategy between each query and the task-specific keys. With queries, keys and the matching strategy in place, the model can evaluate the similarity between the query and keys, selecting the most consistent prompt for the sample to make the final prediction.
4.2. Local Matching Mechanism
@@
Given a sample $\mathbf{x}$, we use the $t$-th e-prompt and g-prompt in combination with ViT for feature extraction to obtain $f_{r}(P_{g};P_{t};\mathbf{x}_{e})[0]$, which serves as the $t$-th query $Q_{t}\in\mathbb{R}^{D}$. The queries generated by using different e-prompts form a query pool $\mathbf{Q}=\{Q_{1},Q_{2},\cdots,Q_{M}\}$. We refer to task-level queries for breadth search as MQ. Next, $Q_{t}$ is locally matched with $\mathbf{k}_{tj}$ by calculating the cosine similarity $\cos\langle Q_{t},\mathbf{k}_{tj}\rangle$. Since different queries encode knowledge from different tasks, enforcing local matching between each query and its corresponding task allows for more effective utilization of task-specific information and helps mitigate cross-query interference. At this point, each prompt has matching scores for all the categories within the task. We need to select the top-$K$ categories with the highest score within a task for aggregation by:
(4) $S^{t}=\underset{\{c_{m}\}_{m=1}^{K}\subseteq[1,|\mathcal{Y}^{t}|]}{\max}\sum_{m=1}^{K}\cos\langle Q_{t},\mathbf{k}_{tc_{m}}\rangle,$
where $S^{t}$ is the aggregated matching score between the sample and the $t$-th prompt. Finally, we select the highest matching score, and get the corresponding index as follows:
(5) $I=\underset{t\subseteq[1,M]}{\operatorname{argmax}}\ S^{t}.$
@@
4.3. Optimization Objective
The overall optimization objective consists of two components: (1) performing classification conditioned on the selected prompt, and (2) selecting prompts whose distribution aligns closely with that of the input sample. Based on the g-prompt and e-prompt, the model is optimized by minimizing the cross-entropy loss between the predicted output and the ground truth label by:
(6) $\mathcal{L}_{\text{cls}}=\text{CE}(W^{T}f_{r}(P_{g};P_{t};\mathbf{x}_{e})[0],y),$
@@ 4.3. Optimization Objective
Since the queries are already guided by the cross-entropy loss in the classification objective, the matching objective focuses solely on aligning the keys with the queries. Hence, the loss for query-key matching can be written as:
(7) $\mathcal{L}_{\text{mch}}=\sum_{i=1}^{M}\mathbb{I}(i=t)\sum_{j=1}^{|\mathcal{Y}^{i}|}(1-\cos\langle Q_{i},\mathbf{k}_{ij}\rangle)\mathbb{I}(j=y),$
where $\mathbb{I}(\cdot)$ is the indicator function. The overall loss function can be written as:
(8) $\mathcal{L}=\mathcal{L}_{\text{cls}}+\mathcal{L}_{\text{mch}}.$
@@
Our MK learns by aligning with the MQ that has already been dispersed through cross-entropy. In contrast, the MK in CPrompt learns by dispersing itself through cross-entropy, since SQ is fixed and not dispersed. Our MK is a component specifically adapted for MQ.
@@
As shown in eq. 7, keys corresponding to other categories and queries corresponding to other prompt are not involved in the loss calculation. This means that during training, only the designated query is needed, indicating that the MQ mechanism not only avoids introducing additional computational overhead, but actually requires one fewer query without a prompt compared to SQSK.
4.4. Efficient Inference of MQMK
MQ integrates knowledge from all tasks for precise querying but comes with a higher inference cost. We aim to design a method that approximates MQ’s query accuracy while achieving SQ’s query speed, seeking a trade-off between accuracy and speed. To ensure a speed comparable to SQ, we maintain a non-expanding single-query architecture. However, to incorporate knowledge from all tasks within the query, we need to enhance the query itself. Here, we design a simple yet effective knowledge fusion method to enhance the query by integrating fused knowledge of all tasks. Specifically, we first perform knowledge fusion by averaging all existing e-prompts in the prompt pool and then obtain the enhanced query using the fused prompt and the g-prompt by:
(9) $Q^{+}=f_{r}(P_{g};\bar{P};\mathbf{x}_{e})[0],$
where $Q^{+}\in\mathbb{R}^{D}$ is the enhanced query and $\bar{P}=\frac{1}{t}\sum_{i=1}^{t}P_{i}$ is the fused prompt. Finally, we replace $Q^{t}$ in MQMK with $Q^{+}$ and select prompts following the same approach as in eq. 4. The above describes the Efficient Inference method of MQMK (MQMK-EI). MQMK-EI is an approximation of MQMK designed to accelerate inference. During training, MQMK is more accurate and more efficient. Therefore, MQMK-EI follows the same training process as MQMK, as described in section 4.3.
4.5. The Perspective of Prompt Ensemble
MQ can be regarded as a prompt ensemble (Lester et al., 2021; Schick and Schütze, 2020) query for multiple tasks in continual learning. Unlike prompt ensemble methods that learn different perspectives of knowledge within a single task, the prompt pool in continual learning stores distinct knowledge across tasks. In the context of continual learning, directly combining or voting over the logits from different prompts may lead to suboptimal performance. MQMK employs a multi-query mechanism in which each query is locally aligned with a specific task. This design helps reduce cross-task interference and knowledge entanglement, making it more suitable for continual learning scenarios characterized by large task discrepancies.
Similar to prompt ensemble methods, we do not need to forward $M$ times through the ViT backbone to obtain the query pool. Instead, we replicate a sample $M$ times, concatenate it with different prompts, and process them in a single batch for parallel inference. If computational resources are highly limited (e.g., the batch size is restricted to 1), MQMK-EI is more efficient. If computational resources are abundant, the computation time of MQMK, when parallelized, is reduced to half compared to the two-stage serial computations of MQMK-EI and SQSK. The MQMK series is training-efficient, achieving approximately a 1/3 speedup compared to the two-stage query process of SQ-based methods. This is a theoretical analysis, and the practical results, the parameters analysis and computational costs during the training phase, are discussed in section 3 of supplementary materials.
@@ 4.5. The Perspective of Prompt Ensemble
CPrompt 78.42 $\pm$ 0.14 5.15 $\pm$ 0.58 77.14 $\pm$ 0.11 5.97 $\pm$ 0.68 74.79 $\pm$ 0.28 7.34 $\pm$ 0.65
MQMK 79.61 $\pm$ 0.27 1.59 $\pm$ 0.26 78.36 $\pm$ 0.35 2.24 $\pm$ 0.20 76.10 $\pm$ 0.17 3.33 $\pm$ 0.26
MQMK-EI 78.50 $\pm$ 0.21 1.24 $\pm$ 0.18 78.00 $\pm$ 0.12 1.20 $\pm$ 0.25 75.82 $\pm$ 0.18 1.87 $\pm$ 0.20
Table 1. Results (%) on Split ImageNet-R under 5-task, 10-task and 20-task settings. Best results are marked in bold. Second-best results are underlined. All our results are over 5 trials. *: Results from (Smith et al., 2023). $\ddagger$}: Results from (Gao et al., 2024). $\dagger$}: Results from (Kurniawan et al., 2024).
Method $A_{T}$ ($\uparrow$) $F_{T}$ $(\downarrow$)
@@
MQMK 91.73 $\pm$ 0.18 2.67 $\pm$ 0.17
MQMK-EI 92.00 $\pm$ 0.29 1.58 $\pm$ 0.18
Table 2. Results (%) on Split CIFAR-100 under 10-task setting. Best results are marked in bold. Second-best results are underlined. All our results are over 5 trials. *: Results from (Smith et al., 2023). $\ddagger$}: Results from (Gao et al., 2024). $\dagger$}: Results from (Kurniawan et al., 2024).
5. Experiments
5.1. Evaluation Benchmarks
Dataset. We shuffle the classes and perform task splitting on three widely used visual datasets for prompt-based continual learning: CIFAR-100 (Krizhevsky et al., 2009), ImageNet-R (Hendrycks et al., 2021), and DomainNet (Peng et al., 2019), in order to align with the problem setup and conduct comprehensive experiments. We divide CIFAR-100 and DomainNet into 10 tasks, while ImageNet is divided into three cases: 5 tasks, 10 tasks, and 20 tasks. The detailed dataset introduction is in section 8 in supplementary materials. The data distributions of DomainNet and ImageNet-R differ significantly from the ImageNet dataset used for model pre-training. This requires the model to continually learn new knowledge from these datasets, rather than relying on the knowledge learned from the pre-training dataset.
Evaluation Metrics. Average accuracy (Kim and Han, 2023; Yan et al., 2021; Li and Hoiem, 2017) and forgetting rate (Chaudhry et al., 2018; Lopez-Paz and Ranzato, 2017) are the two core metrics we use. $A_{T}$ is the average accuracy on tasks 1 to $T$ after the model has learned task $T$, and can be computed by:
(10) $A_{T}=\frac{1}{T}\sum_{t=1}^{T}\text{Accuracy}(t,T),$
@@ 5.1. Evaluation Benchmarks
(11) $F_{T}=\frac{1}{T}\sum_{t=1}^{T}\left(\text{Accuracy}\left(t,t\right)-\text{Accuracy}\left(t,T\right)\right).$
@@
Implementation Details. ViT-B/16 (Dosovitskiy, 2020) pre-trained on ImageNet-21K (Deng et al., 2009) and fine-tuned on ImageNet-1K (Russakovsky et al., 2015) is the backbone we use, so $S$ is set to 16. G-prompt is used in the first two layers of MSA, with a length of 5. The depth and length of e-prompt are discussed in section 5.5. In all experiments, $M$ is set equal to $T$, meaning that only one prompt is learned for each task. Each prompt selects only the category with the highest similarity for aggregation which implies that $K$ is set to 1. For both SQSK and MQMK, a learning rate of 0.005, a batch size of 64, and the Adam (Kingma, 2014) optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ are used in all experiments.
Method $A_{T}$ ($\uparrow$) $F_{T}$ $(\downarrow$)
@@
MQMK 85.62 $\pm$ 0.33 5.51 $\pm$ 0.22
MQMK-EI 84.54 $\pm$ 0.36 3.33 $\pm$ 0.40
Table 3. Results (%) on Split DomainNet under 10-task setting. Best results are marked in bold. Second-best results are underlined. All our results are over 5 trials. Except for MQMK, all the other results come from (Gao et al., 2024).
Competitors and Joint Training. We compare our method with SOTA prompt-based continual learning models, including L2P (Wang et al., 2022b), DualPrompt (Wang et al., 2022a), ESN (Wang et al., 2023a), CODA-P (Smith et al., 2023), EvoPrompt (Kurniawan et al., 2024) and CPrompt (Gao et al., 2024). To ensure a fair comparison, we use SQSK as a baseline in some experiments. SQSK and MQMK have identical prompts, with the only difference being the matching mechanism. SQSK can be seen as a version of DualPrompt with adjusted hyper-parameters, and it has nearly the same performance as DualPrompt reported in tables 1, 2 and 3. Additionally, we jointly train a model using data from all tasks with linear prototypes and prompt tuning as a reference for the upper bound.
5.2. Performance Comparison
As shown in table 1, in the three task settings on ImageNet-R, MQMK achieves SOTA performance. In the 5-task and 10-task splits, MQMK shows no significant performance gap compared to joint training. For the small task split, the continual learning performance with MQMK no longer shows a significant gap compared to traditional joint training with all data. This indicates that the performance on ImageNet-R with the 5-task and 10-task splits is relatively close to its oracle. However, on the 20-task split, MQMK still shows a noticeable performance gap compared to joint training. A possible explanation is that as the number of tasks increases, the limited data per task becomes inadequate for training high-quality e-prompts, ultimately resulting in a degradation of overall performance. Additionally, compared to the high-performing CPrompt and EvoPrompt, MQMK has a clear advantage in terms of forgetting rate. This indicates that MQMK not only achieves higher final performance, but also experiences smaller performance degradation. CODA-P has a low forgetting rate, yet its average accuracy is also relatively low. This suggests that CODA-P exhibits overall stability in performance, but its final performance is average.
Overall, MQMK-EI slightly lags behind MQMK but still achieves SOTA performance. On CIFAR-100 and the 20-task split of ImageNet-R, MQMK-EI achieves performance that is very close to MQMK. This is because CIFAR-100 is a relatively simple dataset with consistent image styles, where knowledge fusion among different prompts introduces minimal loss or interference. In the 20-task split of ImageNet-R, the limited number of samples per task makes it difficult to train high-quality e-prompts. In such cases, knowledge fusion helps to enhance prompt quality to a certain extent.
@@ 5.2. Performance Comparison
Figure 4. Ablation experiments on the 10-task Split ImageNet-R. The average accuracy and matching rate change with the learning process.
Number $A_{T}$ ($\uparrow$) Matching Rate $(\uparrow$)
@@
Table 4. $A_{T}$ (%) and matching rate (%) under keys of different granularities on 10-task Split Imagenet-R.
@@
Figure 5. Average accuracy of MQMK and SQSK on the 10-task Split ImageNet-R under different depths and lengths of the e-prompt. (a) The length is fixed at 80, and the depth varies. (b) The depth is fixed at 10, and the length varies.
5.3. Ablation Study
The components of our method include Multiple Queries (MQ) and Multiple Keys (MK). The corresponding components in existing methods are Single Query (SQ) and Single Key (SK). We discuss the average accuracy and matching rate under four combinations: MQMK, MQSK, SQMK, and SQSK, as well as the inference-efficient variant MQMK-EI, as illustrated in fig. 4. Using SQSK as the baseline and adding MK, $A_{T}$ improves by 2.84% and the matching rate increases by 11.32%. This demonstrates the effectiveness of setting class-level keys for each prompt. By adding MQ, $A_{T}$ decreases by 8.13% and the matching rate drops by 8.92%. This is because, under the supervision of cross-entropy, the features of samples from the same class are more clustered, while the features of samples from different classes are more dispersed. The features of multiple classes in a task have already been dispersed, and in this case, SK can no longer represent the feature distribution of all training samples in a task. When MQ is combined with MK, $A_{T}$ increases by 8.21% and the matching rate improves by 32.82%. This demonstrates that incorporating task-related knowledge significantly enhances retrieval accuracy. In fact, in query-key matching, the improvement of query quality and key quality are complementary. Enhancing both simultaneously has a significant impact on matching rate. MQMK-EI achieves performance comparable to MQMK while incurring a lower query cost.
5.4. Setting Fine-grained Keys
MQSK leads to a performance decline, whereas MQMK improves performance. This highlights the critical importance of learning reliable keys for distribution representation. Based on MQ, we gradually increase the number of keys from 1 to the number of classes in each task (20 class per task in 10-task Split Imagenet-R), and we report their performance in table 4. We divide each task into several class groups based on the number of keys, with each group containing multiple classes. From this perspective, both SK and MK are special cases of $N$-keys.
Method DomainNet ImageNet-R CIFAR-100
L2P 72.31 42.82 48.73
ESN 72.84 70.33 78.50
CPrompt 75.36 64.00 69.37
SQSK 70.55 45.15 49.14
MQMK 82.24 77.97 82.28
MQMK-EI 76.67 75.40 86.17
Table 5. Matching rate (%) of different methods on three 10-task split benchmarks.
It can be observed that as the number of keys increases, both the matching rate and accuracy improve, and the optimal performance is achieved when the key reaches the class level. Since the smallest granularity of our labels is at the class level, when the number of keys exceeds the number of classes, it becomes difficult to provide appropriate supervisory signals for the keys.